The value of π/2 ∫ ((cos 3x+1)/(2 cos x -1)) dx, is 0 A) 2 B) 1 C) 1/2 D) 0

The value of
π/2
  ∫   ((cos 3x+1)/(2 cos x -1)) dx, is
 0
A) 2
B) 1
C) 1/2
D) 0

Grade:

1 Answers

Aditya Gupta
5 years ago
cos3x= 4c^3-3c, where c is cosx
So we have to integrate
(4c^3-3c+1)/(2c-1)= 2c^2+c-1= cosx+cos2x
We can easily integrate this from 0 to pi/2 to obtain the final answer as 1.